OpenStudy (anonymous):

What is the solution of 8 = x / 4

OpenStudy (anonymous):

You're dividing x by 4. So multiply both sides of the equation by 4, hence isolating x and finding its value. \[\large 8 = \frac{ x }{ 4 }\]\[\large 8 \times 4 = \frac{ x }{ 4} \times 4\]\[\large 8 \times 4 = x\] So what is 8 times 4?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Good, so x = 32. (:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw